OpenStudy (anonymous):

A carpenter makes legs for 3-legged stools and 4-legged tables. He makes a total of 57 legs. He could have made legs for how many 3-legged stools?

OpenStudy (anonymous):

@zarax

OpenStudy (anonymous):

@UsArmy3947

OpenStudy (usarmy3947):

4 legged tables : x 3 legged stools : y x + y = 20 y= 20 - x 4x + 3y = 68 Replace y by 20 - x so we only have one letter 4x + 3(20 - x) =68 4x + 60 - 3x =68 x + 60 = 68 x = 8 To find y replace x by 8 y = 20 - 8 y = 12 So you have 8 four legged chairs and 12 three legged stools. Hope I helped!

OpenStudy (anonymous):

?

OpenStudy (usarmy3947):

what are u people talk about i want my medal now

OpenStudy (anonymous):

where does the 20 come from?

OpenStudy (anonymous):

I don't understand this

OpenStudy (anonymous):

me neither

OpenStudy (usarmy3947):

ok i got this from yahoo answers they gave the guy as best response

OpenStudy (anonymous):

He just wants someone 2 give him a medal

OpenStudy (anonymous):

as far as i can see, he could have made \(57\div 3=19\) three legged stools

OpenStudy (usarmy3947):

LIES what is that garbage

OpenStudy (anonymous):

there is no other condition for this problem it does not say how many pieces of furniture were made it only asks, "if he made 57 legs, how many 3 legged stools could he have made?" and the answer is, if he made all stools, there would by 19 of them

OpenStudy (anonymous):

@UsArmy3947 you got to read the question before you google, and read the answer once you have found something to see if it makes sense

OpenStudy (anonymous):

there is no 19

OpenStudy (anonymous):

57 is not divisible by 3?!

OpenStudy (anonymous):

these are the answer choices A. 8 B. 9 C. 10 D. 11

OpenStudy (usarmy3947):

hey who deleted that?

OpenStudy (anonymous):

then there is something missing in the question

OpenStudy (anonymous):

A carpenter makes legs for 3-legged stools and 4-legged tables. He makes a total of 57 legs. He could have made legs for how many 3-legged stools?

OpenStudy (anonymous):

Ops, sorry. Let number of the 3-legged thing is x and 4-legged thing is y. leg numbers: 3*x+4*y = 57 So we are looking for the maximum value of x. Try this: 57 divisible by 3 but he made legs for both 3-legged and 4-legged things so he cant use all of this for 3-legged thing, so subtract 4, it is 53 53 is not divisible by 3, so subtract 4, it is 49 49 is not divisible by 3, su subtract 4, it is 45 45 is divisible by 3 and it is 15. He can make 15 3-legged something.

OpenStudy (anonymous):

That is correct answer, i think

OpenStudy (anonymous):

is this the entire question A carpenter makes legs for 3-legged stools and 4-legged tables. He makes a total of 57 legs. ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

A carpenter makes legs for 3-legged stools and 4-legged tables. He makes a total of 57 legs. He could have made legs for how many 3-legged stools? A. 8 B. 9 C. 10 D. 11

OpenStudy (anonymous):

then something is missing maybe the total number of pieces he made

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

oooh i see maybe what it is asking, it is giving you choices and you have to pick the one that is possible!

OpenStudy (anonymous):

ok lets try 8 if he made 8 stools then he made \(3\times 8=24\) legs for stools that leaves \(57-24=33\) legs for tables

OpenStudy (anonymous):

but that cannot be right because the number of legs for tables must be divisible by 4

OpenStudy (anonymous):

UsArmy! Who is ur gf?

OpenStudy (anonymous):

and 33 is not lets try again with the next number

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\(9\times 3=27\) and \(57-27=30\) but 4 does not go in to 30 evenly so that one is out too

OpenStudy (anonymous):

ok

OpenStudy (usarmy3947):

hey who is the punk that keeps deleting it?

OpenStudy (anonymous):

10 does not work either because \(3\times 10=30\) and \(57-30=27\) four does not go in to 27

OpenStudy (usarmy3947):

show ur face u coward

OpenStudy (usarmy3947):

i bet he is in here

OpenStudy (anonymous):

and finally we try 11

OpenStudy (anonymous):

oh so it's 11?

OpenStudy (anonymous):

\(3\times 11=33\) and \(57-33=24\) which IS divisible by 4 so it is 11

OpenStudy (usarmy3947):

hey now the punk gave me a warning

OpenStudy (anonymous):

oh ok thanks.

OpenStudy (usarmy3947):

if hes here lets go to chat

OpenStudy (anonymous):

Yay it's11 I told u Satellite73 lol

OpenStudy (anonymous):

we needed to understand the question without the answer choices it did not make much sense

OpenStudy (anonymous):

I got it correct. thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

lol satellite u delete it